OpenStudy (anonymous):

Solve the following equation 3-n = 3 ________ __ 2 4

OpenStudy (anonymous):

3/2

Parth (parthkohli):

Multiply both sides by 4. \( \color{Black}{\Rightarrow \Large {4(3- n) \over 2} = 3}\) Multiply both sides by 2. \( \color{Black}{\Rightarrow \Large {4(3 - n) = 6}}\) \( \color{Black}{\Rightarrow 12 - 4n = 6}\) \( \color{Black}{\Rightarrow -4n = -6}\) \( \color{Black}{\Rightarrow 4n = 6}\) \( \color{Black}{\Rightarrow n = 6 \div 4 = 1.5}\)

OpenStudy (anonymous):

Cross Multiply: 4(3 - n) = 3*2 12 - 4n = 6 4n = 12 - 6 n = 6/4 = 3/2..